OpenStudy (dcomer):

1/7(14-7 p) - 2=-2(1/2p+3)+6 solve the equation

OpenStudy (anonymous):

\[\frac{ 1 }{ 7 }(14-7p)-2=-2(\frac{ 1 }{ 2 }p+3)+6\]

OpenStudy (anonymous):

is that your equation?

OpenStudy (dcomer):

yes it is.she asked for us to solve the equation and tell whether is an IDENTITY or NO SOLUTION...:)

OpenStudy (anonymous):

there is definetly a solution to this, so do you know how to start it off?

OpenStudy (anonymous):

lets go ahead and simplify the left side first.\[\frac{ 1 }{ 7 }(14-7p)-2\] \[\frac{ 1 }{ 7 }(14)-\frac{ 1 }{ 7 }(-7p)\]

OpenStudy (anonymous):

do you feel comfortable when im on this step, because i am simply just expanding it.

OpenStudy (dcomer):

yes I feel comfortables

OpenStudy (anonymous):

now how would you complete the expanding, would you mind showing me the simplified form?

OpenStudy (anonymous):